# Invariant Subspaces of Perturbed Backward Shift

### v1

@@ 6. Nearly invariant subspaces
The aim of this section is to establish a connection between our previous results on finite rank perturbations and the concept of nearly invariant subspaces. At first, we set up the notations. Let $\{z_{k}\}_{k=1}^{n}$ be $n$-points (repetition is allowed) in $\mathbb{D}$. Assume that $z_{j}=0$ for some $j=1,\ldots,n$. Define the corresponding Blaschke product $B_{n}$ by
$B_{n}(z)=\prod_{k=1}^{n}\dfrac{z_{k}-z}{1-\bar{z}_{k}z},$
for all $z\in\mathbb{D}$. Clearly, $B_{n}\in H^{\infty}$ is an inner function, and $B_{n}(0)=0$. Then the corresponding model space $\mathcal{K}_{B_{n}}=H^{2}/B_{n}H^{2}$ is $n$-dimensional. Throughout this section, we will fix the Blaschke product mentioned above. We recall the following from Definition 1.4: A closed subspace $\mathcal{M}\subseteq H^{2}$ is nearly $T^{*}_{z,B_{n}}$-invariant if $T_{z}^{*}f\in\mathcal{M}$ for all
@@
for all $z\in\mathbb{D}$. We know that the set of kernel functions $\{k_{w}:w\in\mathbb{D}\}$ is total in $H^{2}$ and also satisfy the reproducing property
@@
for all $f\in H^{2}$ and $w\in\mathbb{D}$. We need the following lemma, which involves Sarason-type perturbations:
Lemma 6.1.
Let $\mathcal{M}$ be a closed subspace of $H^{2}$. If $\mathcal{M}$ is invariant under $T_{z}^{*}-\sum_{i=1}^{n}T_{z}^{*}k_{z_{i}}\otimes k_{z_{i}}$, then
$T_{z}^{*}(\mathcal{M}\cap B_{n}H^{2})\subseteq\mathcal{M}.$
Proof.
Let $f\in\mathcal{M}\cap B_{n}H^{2}$. There exists $g\in H^{2}$ such that $f=B_{n}g$. Then
$(T_{z}^{*}k_{z_{i}}\otimes k_{z_{i}})f=\langle f,k_{z_{i}}\rangle T_{z}^{*}k_{z_{i}}=\langle B_{n}g,k_{z_{i}}\rangle T_{z}^{*}k_{z_{i}}=B_{n}(z_{i})g(z_{i})T_{z}^{*}k_{z_{i}}=0,$
completes the proof of the lemma. ∎
We need another lemma:
@@
$\dim\left(\mathcal{M}\ominus(\mathcal{M}\cap B_{n}H^{2})\right)\leq m.$
@@
Recall that the dimension of the model space $\mathcal{K}_{B_{n}}=H^{2}\ominus B_{n}H^{2}$ is $n$. Suppose $\{f_{1},\ldots,f_{n}\}$ is an orthonormal basis for $\mathcal{K}_{B}$. Set
$\mathcal{L}=span\{P_{\mathcal{M}}f_{1},\ldots,P_{\mathcal{M}}f_{n})\}.$
@@
We are now ready to prove the invariant subspace theorem. It is of importance to point out that the proof follows the same structure as Theorem 2.1. Hence, the subsequent proof will be concise and will exclude details that resemble the proof of Theorem 2.1. Additionally, it is worth recalling that the proof of Theorem 2.1 is partially influenced by the methodology employed by Hitt and Sarason.
@@
Let $\mathcal{M}\subseteq{H}^{2}$ be a nontrivial closed subspace of $H^{2}$. Then $\mathcal{M}$ is nearly $T^{*}_{z,B_{n}}$-invariant if and only if there exist natural numbers $r^{\prime}\leq r\leq n$ and inner function $\Theta\in H^{\infty}_{\mathcal{B}(\mathbb{C}^{r^{\prime}},\mathbb{C}^{r})}$ such that $\Theta(0)=0$ and
@@
where $\{g_{i}\}_{i=1}^{r}$ is an orthonormal basis for $\mathcal{M}\ominus(\mathcal{M}\cap B{H}^{2})$, and
@@
Indeed, if $\mathcal{M}\subseteq B{H}^{2}$, then $\mathcal{M}\subseteq z{H}^{2}$ as $B_{n}(0)=0$. So, for any
@@
we have $a_{0}=0$. By the definition of nearly $T^{*}_{z,B_{n}}$-invariant, we have $T_{z}^{*}f\in\mathcal{M}$, which gives $\sum_{n=1}^{\infty}a_{n}z^{n-1}\in\mathcal{M}\subseteq z{H}^{2}$. But now $a_{1}=0$, and hence, in a similar way, we conclude that $a_{n}=0$ for all $n\geq 0$. Therefore, $f=0$, and hence, =$\mathcal{M}=\{0\}$. This proves the claim, and consequently, we have
@@
$r:=\dim(\mathcal{M}\ominus(\mathcal{M}\cap B_{n}{H}^{2}))\in[1,m].$
@@
with $\|{f}\|^{2}=\|{K}\|^{2}$, where, $K=\sum_{n=0}^{\infty}A_{n}z^{n}\in H^{2}_{\mathbb{C}^{r}}$. With this in mind, define
@@
$T_{z}^{*}f=\frac{G_{0}K}{z}=G_{0}\frac{K-K(0)}{z}=G_{0}(T_{z}^{*}K)\in G_{0}\mathcal{K}_{\Theta}=\mathcal{M},$
from which we conclude that $\mathcal{M}$ is $T^{*}_{z,B_{n}}$-invariant. ∎
Considering the case $B_{1}(z)=z$, we can recover the structure of classical nearly $T_{z}^{*}$-invariant subspaces, which has been described by Hitt [9] and further modified by Sarason [15]. In this case, nearly $T_{z}^{*}$-invariant subspaces are of the form $gK_{\varphi}$ where $\mathcal{K}_{\varphi}\subseteq H^{2}$ is a model space for some inner function $\varphi\in H^{\infty}$ and $g$ is an isometric multiplier on $\mathcal{K}_{\varphi}$ (for instance, see [6]).
@@ 7. Perturbed Toeplitz operators
The purpose of this section is to provide some examples of nearly $T^{*}_{z,B_{n}}$-invariant subspaces. From this perspective, we remind the reader that the kernels of Toeplitz operators are nearly $T_{z}^{*}$-invariant subspaces. Along this line, we prove that the kernels of certain finite-rank perturbations of Toeplitz operators are nearly $T^{*}_{z,B_{n}}$-invariant. Consequently, in addition to the theory of finite rank perturbations of $T_{z}^{*}$, the results of this section can be considered an additional justification for the new notion of nearly $T^{*}_{z,B_{n}}$-invariant subspaces.
@@
where $\mathcal{K}_{B}$ is the model space corresponding to the inner function $B_{n}$. For $\varphi\in L^{\infty}$, the perturbed Toeplitz operator
$T^{\varphi}_{B}:=T_{\varphi}+\sum_{i=1}^{m}\langle\cdot,S^{*}(h_{i})\rangle h_{i},$
is atmost $m$-rank perturbation of the Toeplitz operator $T_{\varphi}$. Given an arbitrary orthonormal basis $\{f_{i}\}_{i=1}^{n}$ of the model space $\mathcal{K}_{B_{n}}$, we will be interested in the finite-rank perturbation of $T_{\varphi}$ as
@@
In the following, we prove that the kernel of the above operator is nearly $T^{*}_{z,B_{n}}$-invariant subspace:
@@
Pick an arbitrary
$f\in\ker(T_{\varphi}+\sum_{i=1}^{n}f_{i}\otimes T_{z}^{*}f_{i})\cap B_{n}H^{2}.$
Our goal is to prove that $T_{z}^{*}f\in\ker(T_{\varphi}+\sum_{i=1}^{n}f_{i}\otimes T_{z}^{*}f_{i})$. Since
@@
and then
$T_{\varphi}f=0.$
Next, we recall that $T_{z}T_{z}^{*}=I-P_{\mathbb{C}}$, where $P_{\mathbb{C}}$ denotes the orthogonal projection of $H^{2}$ onto the one-dimensional space of all constant functions. In the present scenario, $f$ is in $B_{n}H^{2}$ and $B_{n}(0)=0$, which implies $f(0)=0$. Therefore, $P_{\mathbb{C}}f=0$ and hence
@@
Since $f\in BH^{2}$ and $\{f_{i}\}_{i=1}^{n}$ is an orthonormal basis for $\mathcal{K}_{B_{n}}$, it is clear that $\sum_{i=1}^{n}\langle f,f_{i}\rangle f_{i}=0$. Then $T_{\varphi}f=0$ implies that
@@
The converse of the above observation is, of course, not true. For instance, consider $B_{n}$ as $B_{n}=zB_{n-1}$, where $B_{n-1}(0)=0$. Consider the $3$-dimensional subspace $\mathcal{M}$ of $H^{2}$, where
@@
It is simple to see that $\mathcal{M}$ is nearly $T^{*}_{z,B_{n}}$-invariant but neither nearly $T_{z}^{*}$-invariant nor $T_{z}^{*}$-invariant.
In terms of more examples of nearly $T^{*}_{z,B_{n}}$-invariant subspaces, we note that a class of Schmidt subspaces of Hankel operators are nearly $T_{z}^{*}$-invariant subspaces [6]. Given the above illustrations, one would naturally predict that the same would apply to finite-rank perturbations of Hankel operators or perturbations of operators that are related to Hankel operators. It would certainly be fascinating to identify such class operators.

### v2

@@ 6. Nearly invariant subspaces
The aim of this section is to establish a connection between our previous results on finite-rank perturbations and the concept of nearly invariant subspaces. At first, we set up the notations. Let $\{z_{k}\}_{k=1}^{n}$ be $n$-points (repetition is allowed) in $\mathbb{D}$. Assume that $z_{j}=0$ for some $j=1,\ldots,n$. Define
$B_{n}(z)=\prod_{k=1}^{n}\dfrac{z_{k}-z}{1-\bar{z}_{k}z}\qquad(z\in\mathbb{D}),$
the Blaschke product corresponding to $\{z_{k}\}_{k=1}^{n}$. Clearly, $B_{n}\in H^{\infty}$ is an inner function, $B_{n}(0)=0$, and
$\dim\mathcal{K}_{B_{n}}=n.$
Throughout this section, we will fix the Blaschke product mentioned above. We recall the following from Definition 1.4: A closed subspace $\mathcal{M}\subseteq H^{2}$ is nearly $T^{*}_{z,B_{n}}$-invariant if $T_{z}^{*}f\in\mathcal{M}$ for all
@@
for all $z\in\mathbb{D}$. We also know that the set of kernel functions $\{k_{w}:w\in\mathbb{D}\}$ is total in $H^{2}$ and satisfy the reproducing property
@@
for all $f\in H^{2}$ and $w\in\mathbb{D}$. The following general lemma concerns dimension estimations:
@@
$\dim\left(\mathcal{M}\ominus(\mathcal{M}\cap B_{n}H^{2})\right)\leq n.$
@@
Recall that $\dim\mathcal{K}_{B_{n}}=n$. Suppose $\{f_{1},\ldots,f_{n}\}$ is an orthonormal basis for $\mathcal{K}_{B_{n}}$. Set
$\mathcal{L}=span\{P_{\mathcal{M}}f_{1},\ldots,P_{\mathcal{M}}f_{n}\}.$
@@
We are now ready to prove the invariant subspace theorem. The proof follows the same structure as Theorem 2.1. Hence, the subsequent proof will be concise and will exclude details that resemble the proof of Theorem 2.1.
@@
Let $\mathcal{M}\subseteq{H}^{2}$ be a nontrivial closed subspace of $H^{2}$. Then $\mathcal{M}$ is nearly $T^{*}_{z,B_{n}}$-invariant if and only if there exist natural numbers $r^{\prime}\leq r\leq n$, an inner function $\Theta\in H^{\infty}_{\mathcal{B}(\mathbb{C}^{r^{\prime}},\mathbb{C}^{r})}$ with $\Theta(0)=0$, and a unitary operator $\mathcal{J}:\mathcal{K}_{\Theta}\rightarrow\mathcal{M}$ such that
@@
where $\{g_{i}\}_{i=1}^{r}$ is an orthonormal basis for $\mathcal{M}\ominus(\mathcal{M}\cap B_{n}{H}^{2})$, and
@@
Indeed, if $\mathcal{M}\subseteq B_{n}{H}^{2}$, then $\mathcal{M}\subseteq z{H}^{2}$ as $B_{n}(0)=0$. So, for any
@@
we have $a_{0}=0$. By the definition of nearly $T^{*}_{z,B_{n}}$-invariant, we have $T_{z}^{*}f\in\mathcal{M}$, which gives $\sum_{t=1}^{\infty}a_{t}z^{t-1}\in\mathcal{M}\subseteq z{H}^{2}$. But now $a_{1}=0$, and hence, in a similar way, we conclude that $a_{t}=0$ for all $t\geq 0$. Then $f=0$ implying that =$\mathcal{M}=\{0\}$. This proves the claim, and consequently
@@
$r:=\dim(\mathcal{M}\ominus(\mathcal{M}\cap B_{n}{H}^{2}))\in[1,n].$
@@
with $\|{f}\|^{2}=\|{K}\|^{2}$, where, $K=\sum_{t=0}^{\infty}A_{t}z^{t}\in H^{2}_{\mathbb{C}^{r}}$. With this in mind, define
@@
$T_{z}^{*}f=\frac{G_{0}K}{z}=G_{0}\frac{K-K(0)}{z}=G_{0}((T_{z}^{*}\otimes I_{\mathbb{C}^{r}})K)\in G_{0}\mathcal{K}_{\Theta}=\mathcal{M},$
from which we conclude that $\mathcal{M}$ is nearly $T^{*}_{z,B_{n}}$-invariant. ∎
Considering the case $B_{1}(z)=z$, we can recover the structure of classical nearly $T_{z}^{*}$-invariant subspaces, which has been described by Hitt [11] and further modified by Sarason [18]. In this case, nearly $T_{z}^{*}$-invariant subspaces are of the form $gK_{\varphi}$ where $\mathcal{K}_{\varphi}\subseteq H^{2}$ is a model space for some inner function $\varphi\in H^{\infty}$ and $g$ is an isometric multiplier on $\mathcal{K}_{\varphi}$.
Lastly, we establish a connection between invariant subspaces of finite-rank perturbations of $T_{z}^{*}$ and nearly $T^{*}_{z,B_{n}}$-invariant subspaces. This is comparable to the relationship between invariant subspaces of finite-rank perturbations and almost invariant subspaces of $T_{z}^{*}$. The preceding theorem is the underlying basis for the result that follows.
Corollary 6.3.
Let $\mathcal{M}$ be a closed subspace of $H^{2}$. If $\mathcal{M}$ is invariant under $T_{z}^{*}-\sum_{i=1}^{n}T_{z}^{*}k_{z_{i}}\otimes k_{z_{i}}$, then $\mathcal{M}$ is nearly $T^{*}_{z,B_{n}}$-invariant. Conversely, if $\mathcal{M}$ is nearly $T^{*}_{z,B_{n}}$-invariant, then $\mathcal{M}$ is invariant under $T_{z}^{*}-\sum_{i=1}^{r}T_{z}^{*}g_{i}\otimes g_{i}$, where $\{g_{i}\}_{i=1}^{r}$ is an orthonormal basis for $\mathcal{M}\ominus(\mathcal{M}\cap B_{n}{H}^{2})$.
Proof.
Suppose $\mathcal{M}$ is invariant under $T_{z}^{*}-\sum_{i=1}^{n}T_{z}^{*}k_{z_{i}}\otimes k_{z_{i}}$. Let $f\in\mathcal{M}\cap B_{n}H^{2}$. There exists $g\in H^{2}$ such that $f=B_{n}g\in\mathcal{M}$. Then
$(T_{z}^{*}k_{z_{i}}\otimes k_{z_{i}})f=\langle f,k_{z_{i}}\rangle T_{z}^{*}k_{z_{i}}=f(z_{i})T_{z}^{*}k_{z_{i}}=B_{n}(z_{i})g(z_{i})T_{z}^{*}k_{z_{i}}=0,$
completes the proof of the fact that $\mathcal{M}$ is nearly $T^{*}_{z,B_{n}}$-invariant. For the converse direction, assume that $\mathcal{M}$ is nearly $T^{*}_{z,B_{n}}$-invariant. By Theorem 6.2, there exist natural numbers $r^{\prime}\leq r\leq n$, an inner function $\Theta\in H^{\infty}_{\mathcal{B}(\mathbb{C}^{r^{\prime}},\mathbb{C}^{r})}$ with $\Theta(0)=0$, and a unitary operator $\mathcal{J}:\mathcal{K}_{\Theta}\rightarrow\mathcal{M}$ such that
$\mathcal{M}=[g_{1},\ldots,g_{r}]_{1\times r}\mathcal{K}_{\Theta},$
where $\{g_{i}\}_{i=1}^{r}$ is an orthonormal basis for $\mathcal{M}\ominus(\mathcal{M}\cap B_{n}{H}^{2})$, and
$\mathcal{J}(k_{1},\ldots,k_{r})=g_{1}k_{1}+\cdots+g_{r}k_{r},$
for all $(k_{1},\ldots,k_{r})\in\mathcal{K}_{\Theta}$. The remaining part of the proof follows the same line as the converse part of Theorem 3.1. ∎
In a sense, the above result also yields examples of nearly $T^{*}_{z,B_{n}}$-invariant subspaces. We shall see more concrete examples in the following section.
@@ 7. Perturbed Toeplitz operators
The purpose of this section is to provide some examples of nearly $T^{*}_{z,B_{n}}$-invariant subspaces. From this perspective, we remind the reader that the kernels of Toeplitz operators are nearly $T_{z}^{*}$-invariant subspaces [17]. Along this line, we prove that the kernels of certain finite-rank perturbations of Toeplitz operators are nearly $T^{*}_{z,B_{n}}$-invariant. Consequently, in addition to the theory of finite-rank perturbations of $T_{z}^{*}$, the results of this section can be considered an additional justification for the new notion of nearly $T^{*}_{z,B_{n}}$-invariant subspaces.
@@
where $\mathcal{K}_{B_{n}}=H^{2}/B_{n}H^{2}$ is the model space. Given an arbitrary orthonormal basis $\{f_{i}\}_{i=1}^{n}$ of the model space $\mathcal{K}_{B_{n}}$, we will be interested in the finite-rank perturbation of
@@
We now prove that the kernel of this is nearly $T^{*}_{z,B_{n}}$-invariant subspace:
@@
Pick an arbitrary $f\in\ker(T_{\varphi}+\sum_{i=1}^{n}f_{i}\otimes T_{z}^{*}f_{i})\cap B_{n}H^{2}$. Our goal is to prove that $T_{z}^{*}f\in\ker(T_{\varphi}+\sum_{i=1}^{n}f_{i}\otimes T_{z}^{*}f_{i})$. Since
@@
and then $T_{\varphi}f=0$. Next, we recall that $T_{z}T_{z}^{*}=I-P_{\mathbb{C}}$, where $P_{\mathbb{C}}$ denotes the orthogonal projection of $H^{2}$ onto the one-dimensional space of all constant functions. In the present scenario, $f$ is in $B_{n}H^{2}$ and $B_{n}(0)=0$, which implies $f(0)=0$. Therefore, $P_{\mathbb{C}}f=0$ and hence
@@
Since $f\in B_{n}H^{2}$ and $\{f_{i}\}_{i=1}^{n}$ is an orthonormal basis for $\mathcal{K}_{B_{n}}$, it is clear that
$\sum_{i=1}^{n}\langle f,f_{i}\rangle f_{i}=0.$
Then $T_{\varphi}f=0$ implies that
@@
The converse of the above observation is, of course, not true. For instance, consider $B_{n}$ as $B_{n}=zB_{n-1}$, where $B_{n-1}(0)=0$. We consider the $3$-dimensional subspace $\mathcal{M}$ of $H^{2}$, where
@@
It is simple to see that $\mathcal{M}$ is nearly $T^{*}_{z,B_{n}}$-invariant but neither $T_{z}^{*}$-invariant nor nearly $T_{z}^{*}$-invariant.
In terms of more examples of nearly $T^{*}_{z,B_{n}}$-invariant subspaces, we note that a class of Schmidt subspaces of Hankel operators are nearly $T_{z}^{*}$-invariant subspaces [7, 8]. Given the above illustrations, one would naturally predict that the same would apply to finite-rank perturbations of Hankel operators or perturbations of operators that are related to Hankel operators. It would certainly be fascinating to identify such class operators.
In closing, we remark that this paper effectively connects invariant subspaces of three types of operators: finite-rank perturbations of $T_{z}^{*}$, the Blaschke-based backward shift $T_{z,B_{n}}^{*}$, and almost invariant subspaces of $T_{z}^{*}$. We have enhanced the nearly invariant subspace techniques of Hayashi [10], Hitt [11], and Sarason [18]. This enhancement allowed us to tackle the fundamentally challenging invariant subspace problem for perturbed backward shift operators. One should naturally expect further developments along these lines.